Solve the recurrence relationa_n=6a_(n-1)-12a_(n-2)+8a_(n-3) with a_0=-5,a_1=4,and a_2=88.
1
Expert's answer
2020-12-17T19:43:04-0500

characteristic polynomial: x4 - 8x2+16 = 0 roots is 2 and -2 =>

general form of the solutions is an = 2n * (p*n + q) + (-2)n * (r*n + s)

if we know the values of a0 , a1 , a2 , a3 then we get the values p , q , r , s

p = (- 8a0 - 4a1 + 2a2 + a3) / 32 , q = (16a0 + 12 a1 - a3) / 32 ,

r = (- 8a0 + 4a1 + 2a2 - a3) / 32 , s = (16a0 -12a1 + a3) / 32